Two containers, A and B, are separated by a tap. A has a volume of 15 litres and contains one mole
of helium at 25°C. B has a volume of 45 litres and is completely evacuated.
a. Calculate the change in entropy that occurs when the tap is opened.
1
Expert's answer
2020-10-29T05:37:50-0400

"\Delta" S = n*R*ln(V2/V1) = 1*8.314*ln(60/15) = 11.5 J/(mol*K)